OpenStudy (anonymous):

can someone explain to me how to factor each expression without x-boxing it I'd like to learn another faster way I have 72 problems and all I know is x-box???????????

OpenStudy (anonymous):

@Hero @Lena772 @Luigi0210 @tHe_FiZiCx99 @mathmale @Mel98

OpenStudy (lena772):

http://www.mathsisfun.com/algebra/factoring.html check this

OpenStudy (the_fizicx99):

It's a faster method. Need alot of paper though. Looks like a Christmas tree lol. http://amby.com/educate/math/images/factor-tree_96.gif

OpenStudy (anonymous):

lol, thanks guys

OpenStudy (anonymous):

what the monkey is "x box" other than the obvious xmas present?

hero (hero):

For any quadratic polynomial of the form \(ax^2 + bx + c\), Find two numbers that multiply to get \(ac\) yet add to get \(b\) For example: \(x^2 - 16x - 36\) Notice that \(ac = -36\) and \(b = -16\). We need to find two numbers that multiply to get \(-36\) yet add to get \(-16\). The only two numbers that work are \(-18\) and \(2\) since \(-18 \times 2 = -36\) and \(-18 + 2 = -16\) Knowing this we can factor as follows: \(x^2 - 16x - 36\) \(x^2 -18x + 2x - 36\) \(x(x - 18) + 2(x - 18)\) \((x - 18)(x + 2)\)